OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 13x - 6y = 20 7x + 4y = 18 2 1 -2 -1

OpenStudy (nurali):

13x - 6y = 20 7x + 4y = 18 Since we want to know solve for y, we need to get the x terms to cancel out. Notice that the x term in the first equation has a coefficient of 13, and the x term in the second equation has a coefficient of 7. If we multiply each equation by the opposite coefficient, we can get the x terms to cancel out. So multiply the first equation by 7 and the second equation by 13. (7)(13x - 6y) = (7)(20) (13)(7x + 4y) = (13)(18) 91x - 42y = 140 91x + 52y = 234 Notice how the x terms have the same coefficients now? To get them to cancel, subtract bottom equation from top equation. Let us do this one term at a time. We start with the first terms. 91x - 91x = 0 Next do the second terms -42y - 52y = -94y Finally, do the numbers on the right side of the equal sign. 140 - 234 = -94 Now combine all three terms. 0 - 94y = -94 -94y = -94 divide both sides by (-94) -94y/(-94) = -94/(-94) y = 1